OpenStudy (anonymous):

perform the indicated operation and state the domain. given: g(x)=2x+3 h(x)=-x-5 g(x)+h(x)

sam (.sam.):

g(x)=2x+3 h(x)=-x-5 g(x)+h(x) Just substitute, (2x+3)+(-x-5) Solve

OpenStudy (anonymous):

how do I find the domain?

sam (.sam.):

(2x+3)+(-x-5) x-2 graph x-2  Domain =all real numbers